A representative sample of 190 students resulted in a regression equation between y = left hand spans (cm) and x = right hand spans (cm). The least squares regression equation is y = 1.6 + 0.94 x. For a student with a right and left hand span of 28 cm, what is the value of the residual? Give your answer to 2 decimal places.
Answer #1

Regression line given as:

y 1.6 + 0.942. (1)

Given:

Observed y - value = yO = 28 (2)

Predicted y value = yP is obtained by putting x = 28 in (1) as follows:

1.6 + (0.94 × 28) = 27.92 (3)

Residual = Observed y - value - Predicted y - value

= yO - yP

= 28 - 27.92

= 0.08

So,

Answer is:

0.08